OpenStudy (anonymous):

Please I only want help, no answers. The elevation of city A is 96 feet below sea level; the elevation of city C is 24 feet below sea level. The elevation of city D is 1/4 of the elevation of city C. Part A: Write an expression to represent the elevation, in feet, of city D. Part B: Solve the expression and explain why the answer is negative or positive.

OpenStudy (anonymous):

@pooja195?

OpenStudy (anonymous):

@Preetha? @dan815?

OpenStudy (anonymous):

@jlrutherf?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

well do you do 24 - 1/4?

OpenStudy (skullpatrol):

Nope, first you need to assign values for each city's elevation.

OpenStudy (anonymous):

What do you mean by values?

OpenStudy (skullpatrol):

What does it mean for a city to be 24 feet below sea level?

OpenStudy (anonymous):

-24

OpenStudy (skullpatrol):

Very good :-)

OpenStudy (anonymous):

and so A is -94 C is -24 so is D -1/4?

OpenStudy (skullpatrol):

D is not -1/4 because that would be 0.25 feet below sea level, right?

OpenStudy (anonymous):

...idk but i disagree lol

OpenStudy (skullpatrol):

1/4 = 0.25

OpenStudy (anonymous):

I still dont know why it wouldnt be -1/4

OpenStudy (anonymous):

wait....they would all be positive if you divide or multiply right?

OpenStudy (skullpatrol):

Yes, but that is not what the question is asking for.

OpenStudy (anonymous):

ok but it is for part b

OpenStudy (anonymous):

So the expression would be D = -24 - 1/4?

OpenStudy (skullpatrol):

Nope. Let's reread the problem carefully AND think about its meaning

OpenStudy (anonymous):

i did like a 100 times ;(

OpenStudy (skullpatrol):

Please I only want help, no answers. The elevation of city A is 96 feet below sea level = -96; the elevation of city C is 24 feet below sea level = -24. The elevation of city D is 1/4 OF the elevation of city C. Of means multiply. Part A: Write an expression to represent the elevation, in feet, of city D. Part B: Solve the expression and explain why the answer is negative or positive.

OpenStudy (anonymous):

Ok well just to be clear are we adding?

OpenStudy (anonymous):

if we were would it be 24.25?

OpenStudy (skullpatrol):

I think you have your letters mixed up in copying the question down. Pease check them.

OpenStudy (anonymous):

Yea i just checked, i dont

OpenStudy (anonymous):

I will figure this out later. Can you plz check my work on another question? A jar contains 0.85 liter of strawberry juice and 0.50 liter of mango juice. Celina poured 0.15 liter of watermelon juice into the jar. She then drank 0.30 liter of the mixture. Part A: Write an expression to represent the total amount of juice left in the jar. (5 points) Part B: Identify which property is used in each step to simplify the expression This is my work: 0.85 = strawberry 0.50 = mango Added 0.15 of watermelon into the jar. Drank 0.30. Part A: Total = 1.5 Left = 1.2 The expression is J = (S + M) + (W - D) J = (.85 + .50) + ( .15 - .30) J = 1.2 liters Part B: The property being used is Associative.

OpenStudy (mathstudent55):

@beckyg111 This may help you understand what the first problem is asking. John dug a hole 12 feet deep. Bob dug a hole 1/4 the depth of John's hole. How deep is Bob's hole?

OpenStudy (skullpatrol):

The elevation of city A is 96 feet below sea level; the elevation of city C is 24 feet below sea level. The elevation of city D is 1/4 of the elevation of city C. Part A: Write an expression to represent the elevation, in feet, of city D. D = 0.25*C Answer Part B: Solve the expression and explain why the answer is negative or positive. D = 0.25*(-24) = -6 feet below sea level The answer is negative because it is below sea level. Answer

OpenStudy (anonymous):

@mathstudent55 would that be 3?

OpenStudy (anonymous):

@skullpatrol D = 1/16 ?

OpenStudy (skullpatrol):

D = -6

OpenStudy (mathstudent55):

Yes. How did you get 3? You get 3 by multiplying 12 ft by 1/4, which is the same as 12 ft * 0.25 It's the same in your problem. City C has an elevation of 24 ft below sea level. That elevation can be written as a negative number, -24 ft. City D has an elevation 1/4 of the elevation of city C, so city D's elevation is 1/4 * C Notice the operation is 1/4 * C. The elevation of D is not just 1/4. The elevation of city D is not simply -1/4 ft, or -0.25 ft. It is 1/4 * (-24 ft)

OpenStudy (anonymous):

Ok sooo multiplication was what i needed to do?

OpenStudy (mathstudent55):

1/4 * C = 1/4 * (-24 ft) = 0.25 * (-24 ft) = -6 ft The elevation of D is 6 ft below sea level.

OpenStudy (mathstudent55):

Correct. Just like in my example you multiplied 1/4 * 12 ft, here you multiply 1/4 * (-24 ft)

OpenStudy (anonymous):

ok that makes more sense

OpenStudy (skullpatrol):

1/4 OF means "multiply"

OpenStudy (mathstudent55):

If you read @skullpatrol 's responses, you'll see that is what he was trying to tell you.

OpenStudy (anonymous):

lol sorry

OpenStudy (mathstudent55):

Don't be sorry. Keep studying and asking questions. That's how you learn. I like your attitude that you don't want answers; you want to understand how to solve the problem.

OpenStudy (skullpatrol):

What does it mean for a city to be 24 feet below sea level?